# Prompt-guided Disentangled Representation for Action Recognition

### v3

@@ 1 Introduction
Action recognition is one of the fundamental tasks in the field of video understanding [16], and it remains an active topic in the vision research community. Current mainstream methods can be categorized into two types based on their data processing approach. The first type is based on complete video data. These methods typically input the untrimmed video into 2D Convolution Neural Network (CNN), 3D CNN, or Vision Transformer (ViT) [7], using the network to extract features and then performing classification using the extracted features [10, 42, 4, 45, 27, 30, 27]. The second type is object-centric, where object detectors first crop objects from the video and then input them into ViT or Graph Neural Network (GNN) [53, 33, 57, 48, 54, 55, 48, 53]. Some methods [59, 20, 47, 35] also construct Spatio-temporal Scene Graph (SSG) from these objects and then feed them into the network. However, these methods are essentially information extraction processes that handle all actions in the video at once, making it difficult to model the interactions between objects in complex scenes [48, 55]. In this paper, we aim to analyze actions with a disentangled approach. For instance, in a video with multiple actions, we want the network to focus on analyzing some specified action(s). Moreover, we construct SSGs for this video and aim to disentangle the subgraph, which consist of the specified action-relevant nodes, enabling better understanding of the actions (as shown in Figure 1).
@@
Existing Disentangled Representaion Learning (DRL) methods typically approach disentanglement from different perspectives. Some methods focus on attribute disentanglement, aiming to separate content from background [22, 52, 17]. Others adopt an information perspective by disentangling representations into structural invariants and local variations (e.g., shared information across nodes and private, node-specific features) [6, 58, 12, 34]. From a structural perspective, some approaches disentangle graphs into a fixed number components(e.g., latent representation, subgraphs) [26, 25, 24, 6, 58, 12]. A task perspective has also been explored, where complex tasks are divided into multiple subtasks to facilitate more structured representation learning [29]. In action recognition, we argue that disentanglement should be approached from the action perspective, enabling the dynamic separation of semantic factors related to any specified actions. Specifically, we expect the network to analyze any actions we specify and to disentangle the corresponding subgraph from the whole SSG, enabling more targeted and interpretable action understanding.
To this end, in order to enable the model to disentangle representations conditioned on any specified action, we draw inspiration from recent advances in prompt tuning [31, 8, 39, 21, 56, 2] and propose Prompt-guided Disentangled Representation for Action Recognition (ProDA). As depict in Figure 1 (a), ProDA disentangles the SSG into two complementary parts: one corresponding to the specified actions, and the other to the remaining actions. To achieve this, we construct an action-aware disentanglement network that can disentangle any specified actions from SSG based on prompt guidance. Specifically, we introduce the Dynamic Prompt Module (DPM), which generates action-aware prompts by combining the multi-hot vector of specified actions with the SSG features. These prompts guide the Graph Parsing Neural Network (GPNN) to perform action-aware disentanglement. Furthermore, we propose a Video Graph Parsing Neural Network (VGPNN), a GPNN variant tailored for SSG, to perform action disentanglement. To ensure temporal consistency and preserve the differences across SSGs from different videos, we introduce a Video Graph Normalization (VGNorm) module, which is applied after each VGPNN layer. Finally, we introduce an Action Disentanglement Loss (AD Loss) to supervise the separated representations. To ensure that the disentangled features retain essential information, we incorporate a reconstruction term, promoting true representation disentanglement rather than mere feature extraction.
@@ 2 Related works
Existing action recognition methods can be broadly categorized into three types. First type is CNN-based methods. To compensate for the lack of temporal modeling in 2D CNNs, two-stream networks introduce an additional temporal stream (e.g., optical flow)[10, 9, 41, 44, 49]. By jointly modeling spatial and temporal features, 3D CNNs are able to better capture spatio-temporal information and thus yield improved performance [42, 4, 51]. The second type is ViT-based methods. Most methods either treat entire frames as input tokens or divide each frame into patches before feeding them into the model [13, 45, 27, 30, 27]. These methods treat the video as a whole, which makes it difficult to model interactions among objects, and often introduces substantial noise. Our work alleviate this by disentangling actions from the SSG, enabling more precise modeling of object interactions. The third type is Objected-centric methods. Object-centric methods construct more compact and efficient video representations by focusing on regions of interest (e.g., objects in the video) [53], and enabling generalization of unseen interaction relationships through structured representations [53, 33, 57, 48, 59]. Nevertheless, these methods remain limited in their ability to effectively model interactions among objects. The incorporation of SSG [59, 20] effectively addresses the difficulty in modeling object interactions. In scene graphs, videos are represented as spatio-temporal graph structures, with objects serving as nodes (e.g., human, cup) and their interactions or spatial-relationship (e.g., holding, above) modeled as edges. [35] processes the scene graph of each frame using a Graph Neural Network (GNN). On the other hand, [47] flattens the scene graphs and feeds them into a ViT [7], leveraging the long-range modeling capability of self-attention [43] to effectively capture interaction transitions. These SSG-based methods still adopt feature extraction approaches that process all actions in a video, which makes it challenging to model interactions between different objects in multi-action scenarios. In contrast, our method disentangles actions from complex scenes, enabling better understanding of the actions
@@ 3 Methodology
Our method aims to disentangle any specified actions from multi-action scenarios. This disentanglement provides more concise feature representation for analyzing the specified actions, leading to improved action understanding. The overall architecture of our method is illustrated in Figure  2, we first convert the input video into a SSG. This graph is then processed by a video encoder (specifically, Dual-AI [15]) to model temporal and spatial interactions among nodes within the SSG. Next, guided by the DPM, the video features are fed into VGPNN for representation disentanglement. Finally, to further enhance the disentanglement, in addition to standard classification supervision, we introduce both a reconstruction loss and a disentanglement loss as constraints. We generate the SSG [38] using existing methods and adopt a off-the-shelf video encoder [15]; the following focuses on the remaining components of the network architecture illustrated in Figure 2.
@@ 3.2 Dynamic Prompt Module (DPM)
The DPM is a key component of our method. Since action disentanglement is driven by prompts, ambiguous prompts can make it hard for the model to focus on the specified action-relevant regions, leading to failure in disentangling the specified actions. we generate dynamic prompts based on the node features of the SSG, which guide the model to attend to action-relevant nodes and achieve better disentanglement. In this section, we introduce how prompts are generated based on arbitrary action specifications. We first describe the design of action specifications, followed by their construction strategy. Finally, we show how prompts are generated from the specifications, which guide the VGPNN to disentangle actions. Inspired by [8, 31], we inject learnable prompts into SSG feature space to accomplish guidance.
@@ 3.3 Video Graph Parsing Neural Network (VGPNN)
We aim to disentangle subgraphs corresponding to specified actions. To achieve this, we require a network that, under the guidance of prompts, can parse action-relevant subgraphs by directionally propagating action-specific information. Inspired by GPNN [37], we design a variant for SSG, called VGPNN. GPNN parses action-relevant nodes by generating dynamic weights that control the direction of information flow. It consists of three key steps: Link, Message, and Update. In the Link step, it constructs edges between the same object nodes across different frames and learns aggregation weights for each timestamp. In the Message step, it multiplies these weights with messages, where each message is composed of a node’s features and its edge features. In the Update step, it selects nodes from a specific frame as the initial hidden states, and updates them sequentially using the weighted messages from different time steps. However, GPNN has several limitations when applied to SSGs. 1)It only updates a node from a single frame, ignoring joint modeling of nodes across the entire video. 2) It performs shallow temporal modeling, as the learned weights are used only once for aggregation and are not further refined or propagated through deeper temporal structures.
@@
Video Graph Parsing Neural Network (VGPNN). VGPNN is designed to disentangle features by decomposing a SSG into two subgraphs, guided by a given AS. To achieve selective information aggregation, with the same to GPNN [37], each VGPNN layer processes the entire SSG in three sequential steps. In Link step, we perform temporal modeling of edges $e^{j,k}=\{e_{1}^{j,k},e_{i}^{j,k},\dots,e_{N}^{j,k}\}$ via self-attention [43] over the same interaction across time. In Message step, we generate dynamic weights from the edge features, which are then multiplied with the messages, where each message is composed of a node’s features and its edge features. In Update step, we update all nodes by adding the weighted messages to the corresponding nodes directly. A MLP is then used to fuse the updated and original node features. More details please refer to Sec. B.
@@
Video Graph Normalization (VGNorm). SSG exhibits characteristics of both graph structure and spatio-temporal properties of videos. However, many commonly used normalization methods tend to focus on only one aspect. For instance, BatchNorm [18] primarily normalizes across the temporal batch dimension, while LayerNorm [1] operates along the node (spatial) dimension to enforce spatial consistency. This one-sided focus limits network performance. To ensure temporal consistency and preserve the differences across SSGs from different videos, our method focuses on two key aspects. First, we use the frame-wise global mean $\mu_{g}\in\mathbb{R}^{N}$ to ensure temporal consistency across frames. During training, we compute the mean $\mu\in\mathbb{R}^{N}$ for each batch in real-time and update the $\mu_{g}\in\mathbb{R}^{N}$ using a momentum-based approach. Specifically, for batchsize $B$ nodes features $\mathcal{F}^{node}=\{\mathcal{F}^{node}_{1},\mathcal{F}_{2}^{node},\dots,\mathcal{F}_{B}^{node}\}\in\mathbb{R}^{B\times N\times M\times D}$, where $\mathcal{F}_{b}^{node}\in\mathbb{R}^{N\times M\times D}$ is the $b$-th video nodes features, the global mean is computed as:
@@ 3.4 Optimization
where $\mathcal{F}^{u}_{b},\mathcal{F}^{s}_{b}\in\mathbb{R}^{N\times M\times D}$ represent the action representations disentangled by UAP and SAP of $b$-th video, respectively. $\phi_{b},\psi_{b}$ are measurable functions [14]. $\mathtt{Cor}(\cdot),\mathtt{Var}(\cdot),m_{1}$ represent covariance, variance and margin, respectively.
@@
SGFB+NL [20] K400 RGB 44.3 SGFB+NL [20] K400 RGB+SSG 60.3
I3D+NL [4] K400 RGB 37.5 LaIAR [47] K400 RGB+Bbox 63.6
OR2G+NL [35] K400 RGB 44.9 OR2G [35] K400 RGB+SSG 63.3
LaIAR+NL [47] K400 RGB 45.1 LaIAR+NL [47] K400 RGB+Bbox 67.4
VicTR(B/16) [23] CLIP RGB 50.1 OR2G+NL [35] K400 RGB+SSG 67.5
VicTR(L/14) [23] CLIP RGB 57.6 Ours(B/16) CLIP RGB+SSG 71.1
@@
STTran [5] 3.31 71.29 62.91 42.67 22.14 -
HORT  [19] 3.75 78.57 67.78 50.33 26.96 -
SlowFast [11] 5.00 81.66 74.00 52.74 26.82 -
ACARN  [36] 5.44 82.85 75.22 56.53 31.77 -
SportHHI  [50] (baseline) 10.69 89.25 82.93 68.13 43.72 111.18M
@@ 4.1 Experiment Setup
Dataset. (1) The Charades dataset [40] consists of 9,848 videos, each with an average duration of 30 seconds. It covers 157 action categories, with each video containing an average of 6.8 distinct actions. Multiple actions can occur simultaneously, making the recognition task particularly challenging. Based on Charades, the Action Genome dataset [20] provides fine-grained annotations by decomposing actions and focusing on video clips where the actions take place. It contains 234K keyframes, with annotations for 476K object bounding boxes and 1.72 million object relationships. (2) The SportsHHI dataset [50] is a specially designed dataset focusing on human-human interaction in sports. It covers basketball and volleyball games, and video data are selected from MultiSports [28]. This dataset labels 34 interaction classes with 118,075 human bounding boxes and 50,649 interaction instances. Following [47, 50], we conduct evaluations on Charades for multi-label action recognition and on SportsHHI for Human-Human interaction Classification (HHICls), adopting mean Average Precision (mAP) as the metric.
@@ 4.2 Compared with State-of-The Art Methods
Charades. We compare the action recognition accuracy of the proposed method and the State-of-The-Art methods (SoTA) on the Charades. Table 1 presents our results on the Charades dataset. ProDA achieves state-of-the-art performance. Our method outperforms several CNN- and ViT-based approaches, including I3D [4] and VicTR [23]. This demonstrates that our approach can better capture object interactions within structured data (e.g., SSG), enabling a deeper understanding of video. Notably, our method with a CLIP-B/16 [38] still surpasses VicTR (L/14) [23], even though we only use an image encoder, while VicTR [23] additionally employs a text encoder. Furthermore, our method also outperforms several approaches that leverage SSG [35, 47]. This demonstrates that explicitly disentangling and processing actions separately leads to a more accurate understanding of complex actions. It is also worth noting that our method achieves this performance without requiring any additional external models.

### v4

@@ 1 Introduction
Action recognition is one of the fundamental tasks in the field of video understanding [hassner2013critical], and it remains an active topic in the vision research community. Current mainstream methods can be categorized into two types based on their data processing approach. The first type is based on complete video data. These methods typically input the untrimmed video into 2D Convolution Neural Network (CNN), 3D CNN, or Vision Transformer (ViT) [vit], using the network to extract features and then performing classification using the extracted features [feichtenhofer2017spatiotemporal, tran2015learning, carreira2017quo, wang2024omnivid, li2022mvitv2, liu2022video, li2022mvitv2]. The second type is object-centric, where object detectors first crop objects from the video and then input them into ViT or Graph Neural Network (GNN) [zhang2024object, materzynska2020something, zhou2023can, wang2018videos, zhang2020temporal, zhang2019structured, wang2018videos, zhang2024object]. Some methods [zhuo2019explainable, actiongenome, lair, ou2022object] also construct Spatio-temporal Scene Graph (SSG) from these objects and then feed them into the network. However, these methods are essentially information extraction processes that handle all actions in the video at once, making it difficult to model the interactions between objects in complex scenes [wang2018videos, zhang2019structured]. In this paper, we aim to analyze actions with a disentangled approach. For instance, in a video with multiple actions, we want the network to focus on analyzing some specified action(s). Moreover, we construct SSGs for this video and aim to disentangle the subgraph, which consist of the specified action-relevant nodes, enabling better understanding of the actions (as shown in Figure 1).
@@
Existing Disentangled Representaion Learning (DRL) methods typically approach disentanglement from different perspectives. Some methods focus on attribute disentanglement, aiming to separate content from background [PSGAN, DARLING, hsieh2018learning]. Others adopt an information perspective by disentangling representations into structural invariants and local variations (e.g., shared information across nodes and private, node-specific features) [denton2017unsupervised, zhu2020s3vae, gan2025dfdnet, mo2023disentangled]. From a structural perspective, some approaches disentangle graphs into a fixed number components(e.g., latent representation, subgraphs) [graph_dis, li2022disentangled, li2021disentangled, denton2017unsupervised, zhu2020s3vae, gan2025dfdnet]. A task perspective has also been explored, where complex tasks are divided into multiple subtasks to facilitate more structured representation learning [li2024disentangled]. In action recognition, we argue that disentanglement should be approached from the action perspective, enabling the dynamic separation of semantic factors related to any specified actions. Specifically, we expect the network to analyze any actions we specify and to disentangle the corresponding subgraph from the whole SSG, enabling more targeted and interpretable action understanding.
To this end, in order to enable the model to disentangle representations conditioned on any specified action, we draw inspiration from recent advances in prompt tuning [ma2024hetgpt, fang2023universal, shi2023dept, jia2022visual, zhou2022conditional, bahng2022exploring] and propose Prompt-guided Disentangled Representation for Action Recognition (ProDA). As depict in Figure 1 (a), ProDA disentangles the SSG into two complementary parts: one corresponding to the specified actions, and the other to the remaining actions. To achieve this, we construct an action-aware disentanglement network that can disentangle any specified actions from SSG based on prompt guidance. Specifically, we introduce the Dynamic Prompt Module (DPM), which generates action-aware prompts by combining the multi-hot vector of specified actions with the SSG features. These prompts guide the Graph Parsing Neural Network (GPNN) to perform action-aware disentanglement. Furthermore, we propose a Video Graph Parsing Neural Network (VGPNN), a GPNN variant tailored for SSG, to perform action disentanglement. To ensure temporal consistency and preserve the differences across SSGs from different videos, we introduce a Video Graph Normalization (VGNorm) module, which is applied after each VGPNN layer. Finally, we introduce an Action Disentanglement Loss (AD Loss) to supervise the separated representations. To ensure that the disentangled features retain essential information, we incorporate a reconstruction term, promoting true representation disentanglement rather than mere feature extraction.
@@ 2 Related works
Existing action recognition methods can be broadly categorized into three types. First type is CNN-based methods. To compensate for the lack of temporal modeling in 2D CNNs, two-stream networks introduce an additional temporal stream (e.g., optical flow)[feichtenhofer2017spatiotemporal, feichtenhofer2016convolutional, simonyan2014two, wang2013action, wang2017spatiotemporal]. By jointly modeling spatial and temporal features, 3D CNNs are able to better capture spatio-temporal information and thus yield improved performance [tran2015learning, carreira2017quo, xie2018rethinking]. The second type is ViT-based methods. Most methods either treat entire frames as input tokens or divide each frame into patches before feeding them into the model [girdhar2019video, wang2024omnivid, li2022mvitv2, liu2022video, li2022mvitv2]. These methods treat the video as a whole, which makes it difficult to model interactions among objects, and often introduces substantial noise. Our work alleviate this by disentangling actions from the SSG, enabling more precise modeling of object interactions. The third type is Objected-centric methods. Object-centric methods construct more compact and efficient video representations by focusing on regions of interest (e.g., objects in the video) [zhang2024object], and enabling generalization of unseen interaction relationships through structured representations [zhang2024object, materzynska2020something, zhou2023can, wang2018videos, zhuo2019explainable]. Nevertheless, these methods remain limited in their ability to effectively model interactions among objects. The incorporation of SSG [zhuo2019explainable, actiongenome] effectively addresses the difficulty in modeling object interactions. In scene graphs, videos are represented as spatio-temporal graph structures, with objects serving as nodes (e.g., human, cup) and their interactions or spatial-relationship (e.g., holding, above) modeled as edges. [ou2022object] processes the scene graph of each frame using a Graph Neural Network (GNN). On the other hand, [lair] flattens the scene graphs and feeds them into a ViT [vit], leveraging the long-range modeling capability of self-attention [transformer] to effectively capture interaction transitions. These SSG-based methods still adopt feature extraction approaches that process all actions in a video, which makes it challenging to model interactions between different objects in multi-action scenarios. In contrast, our method disentangles actions from complex scenes, enabling better understanding of the actions
@@ 3 Methodology
Our method aims to disentangle any specified actions from multi-action scenarios. This disentanglement provides more concise feature representation for analyzing the specified actions, leading to improved action understanding. The overall architecture of our method is illustrated in Figure  2, we first convert the input video into a SSG. This graph is then processed by a video encoder (specifically, Dual-AI [han2022dual]) to model temporal and spatial interactions among nodes within the SSG. Next, guided by the DPM, the video features are fed into VGPNN for representation disentanglement. Finally, to further enhance the disentanglement, in addition to standard classification supervision, we introduce both a reconstruction loss and a disentanglement loss as constraints. We generate the SSG [clip] using existing methods and adopt a off-the-shelf video encoder [han2022dual]; the following focuses on the remaining components of the network architecture illustrated in Figure 2.
@@ 3.2 Dynamic Prompt Module (DPM)
The DPM is a key component of our method. Since action disentanglement is driven by prompts, ambiguous prompts can make it hard for the model to focus on the specified action-relevant regions, leading to failure in disentangling the specified actions. we generate dynamic prompts based on the node features of the SSG, which guide the model to attend to action-relevant nodes and achieve better disentanglement. In this section, we introduce how prompts are generated based on arbitrary action specifications. We first describe the design of action specifications, followed by their construction strategy. Finally, we show how prompts are generated from the specifications, which guide the VGPNN to disentangle actions. Inspired by [fang2023universal, ma2024hetgpt], we inject learnable prompts into SSG feature space to accomplish guidance.
@@ 3.3 Video Graph Parsing Neural Network (VGPNN)
We aim to disentangle subgraphs corresponding to specified actions. To achieve this, we require a network that, under the guidance of prompts, can parse action-relevant subgraphs by directionally propagating action-specific information. Inspired by GPNN [gpnn], we design a variant for SSG, called VGPNN. GPNN parses action-relevant nodes by generating dynamic weights that control the direction of information flow. It consists of three key steps: Link, Message, and Update. In the Link step, it constructs edges between the same object nodes across different frames and learns aggregation weights for each timestamp. In the Message step, it multiplies these weights with messages, where each message is composed of a node’s features and its edge features. In the Update step, it selects nodes from a specific frame as the initial hidden states, and updates them sequentially using the weighted messages from different time steps. However, GPNN has several limitations when applied to SSGs. 1)It only updates a node from a single frame, ignoring joint modeling of nodes across the entire video. 2) It performs shallow temporal modeling, as the learned weights are used only once for aggregation and are not further refined or propagated through deeper temporal structures.
@@
Video Graph Parsing Neural Network (VGPNN). VGPNN is designed to disentangle features by decomposing a SSG into two subgraphs, guided by a given AS. To achieve selective information aggregation, with the same to GPNN [gpnn], each VGPNN layer processes the entire SSG in three sequential steps. In Link step, we perform temporal modeling of edges $e^{j,k}=\{e_{1}^{j,k},e_{i}^{j,k},\dots,e_{N}^{j,k}\}$ via self-attention [transformer] over the same interaction across time. In Message step, we generate dynamic weights from the edge features, which are then multiplied with the messages, where each message is composed of a node’s features and its edge features. In Update step, we update all nodes by adding the weighted messages to the corresponding nodes directly. A MLP is then used to fuse the updated and original node features. More details please refer to Sec. B.
@@
Video Graph Normalization (VGNorm). SSG exhibits characteristics of both graph structure and spatio-temporal properties of videos. However, many commonly used normalization methods tend to focus on only one aspect. For instance, BatchNorm [batchnorm] primarily normalizes across the temporal batch dimension, while LayerNorm [layernorm] operates along the node (spatial) dimension to enforce spatial consistency. This one-sided focus limits network performance. To ensure temporal consistency and preserve the differences across SSGs from different videos, our method focuses on two key aspects. First, we use the frame-wise global mean $\mu_{g}\in\mathbb{R}^{N}$ to ensure temporal consistency across frames. During training, we compute the mean $\mu\in\mathbb{R}^{N}$ for each batch in real-time and update the $\mu_{g}\in\mathbb{R}^{N}$ using a momentum-based approach. Specifically, for batchsize $B$ nodes features $\mathcal{F}^{node}=\{\mathcal{F}^{node}_{1},\mathcal{F}_{2}^{node},\dots,\mathcal{F}_{B}^{node}\}\in\mathbb{R}^{B\times N\times M\times D}$, where $\mathcal{F}_{b}^{node}\in\mathbb{R}^{N\times M\times D}$ is the $b$-th video nodes features, the global mean is computed as:
@@ 3.4 Optimization
where $\mathcal{F}^{u}_{b},\mathcal{F}^{s}_{b}\in\mathbb{R}^{N\times M\times D}$ represent the action representations disentangled by UAP and SAP of $b$-th video, respectively. $\phi_{b},\psi_{b}$ are measurable functions [gretton2005kernel]. $\mathtt{Cor}(\cdot),\mathtt{Var}(\cdot),m_{1}$ represent covariance, variance and margin, respectively.
@@
SGFB+NL [actiongenome] K400 RGB 44.3 SGFB+NL [actiongenome] K400 RGB+SSG 60.3
I3D+NL [carreira2017quo] K400 RGB 37.5 LaIAR [lair] K400 RGB+Bbox 63.6
OR2G+NL [ou2022object] K400 RGB 44.9 OR2G [ou2022object] K400 RGB+SSG 63.3
LaIAR+NL [lair] K400 RGB 45.1 LaIAR+NL [lair] K400 RGB+Bbox 67.4
VicTR(B/16) [victr] CLIP RGB 50.1 OR2G+NL [ou2022object] K400 RGB+SSG 67.5
VicTR(L/14) [victr] CLIP RGB 57.6 Ours(B/16) CLIP RGB+SSG 71.1
@@
STTran [cong2021spatial] 3.31 71.29 62.91 42.67 22.14 -
HORT  [hort] 3.75 78.57 67.78 50.33 26.96 -
SlowFast [slowfast] 5.00 81.66 74.00 52.74 26.82 -
ACARN  [acarn] 5.44 82.85 75.22 56.53 31.77 -
SportHHI  [sportshhi] (baseline) 10.69 89.25 82.93 68.13 43.72 111.18M
@@ 4.1 Experiment Setup
Dataset. (1) The Charades dataset [charades] consists of 9,848 videos, each with an average duration of 30 seconds. It covers 157 action categories, with each video containing an average of 6.8 distinct actions. Multiple actions can occur simultaneously, making the recognition task particularly challenging. Based on Charades, the Action Genome dataset [actiongenome] provides fine-grained annotations by decomposing actions and focusing on video clips where the actions take place. It contains 234K keyframes, with annotations for 476K object bounding boxes and 1.72 million object relationships. (2) The SportsHHI dataset [sportshhi] is a specially designed dataset focusing on human-human interaction in sports. It covers basketball and volleyball games, and video data are selected from MultiSports [li2021multisports]. This dataset labels 34 interaction classes with 118,075 human bounding boxes and 50,649 interaction instances. Following [lair, sportshhi], we conduct evaluations on Charades for multi-label action recognition and on SportsHHI for Human-Human interaction Classification (HHICls), adopting mean Average Precision (mAP) as the metric.
@@ 4.2 Compared with State-of-The Art Methods
Charades. We compare the action recognition accuracy of the proposed method and the State-of-The-Art methods (SoTA) on the Charades. Table 1 presents our results on the Charades dataset. ProDA achieves state-of-the-art performance. Our method outperforms several CNN- and ViT-based approaches, including I3D [carreira2017quo] and VicTR [victr]. This demonstrates that our approach can better capture object interactions within structured data (e.g., SSG), enabling a deeper understanding of video. Notably, our method with a CLIP-B/16 [clip] still surpasses VicTR (L/14) [victr], even though we only use an image encoder, while VicTR [victr] additionally employs a text encoder. Furthermore, our method also outperforms several approaches that leverage SSG [ou2022object, lair]. This demonstrates that explicitly disentangling and processing actions separately leads to a more accurate understanding of complex actions. It is also worth noting that our method achieves this performance without requiring any additional external models.